OpenStudy (anonymous):

Simplify the following expression. 2x - 2y + 5z - 2x - y + 3z

OpenStudy (anonymous):

-3y+8z

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

@Jiimy20110 It seems you are satisfied with just answers. Full Solution: 2x - 2y + 5z - 2x - y + 3z = 2x - 2x - 2y - y + 5z + 3z = 0 - 3y + 8z = 8z - 3y.